# Linear Complementary Pairs of Quasi-Cyclic and Quasi-Twisted Codes

### v1

@@ Abstract
In this paper, we provide a polynomial characterization of linear complementary pairs of quasi-cyclic and quasi-twisted codes of index $2$. We also give several examples of linear complementary pairs of quasi-cyclic and quasi-twisted codes with (almost) optimal security parameters.
@@ 1 Introduction
In 2018, Carlet et al. [9] characterized LCPs of quasi-cyclic codes using the Chinese Remainder Theorem (CRT) decomposition of codes and also provided necessary and sufficient conditions for pairs of constacyclic and double circulant codes to be an LCP of codes. Moreover, they proved that if $(C,D)$ is an LCP of constacyclic or 2D cyclic code, then $C$ and $D^{\perp}$ are permutation equivalent. Güneri et al. [12] showed the equivalence of $C$ and $D^{\perp}$ holds for an LCP of $mD$-cyclic codes as well. These results show that finding LCD codes and LCPs of codes is the same in the context of security parameters. However, $C$ and $D^{\perp}$ need not be permutation equivalent for an LCP of quasi-cyclic codes [9, Example 3.3]. In this work, we give a polynomial characterization of LCPs of quasi-cyclic and quasi-twisted codes with index $2$. The rest of the paper is organized as follows.
@@ 2 Preliminaries
Note that if $C$ has parameters $[n,k]$, then $D$ must have parameters $[n,n-k]$. Linear complementary dual (LCD) codes are a special class of LCPs of codes where $D$ is the dual code of $C$. Equivalently, $C$ is an LCD code if $C\cap C^{\perp}=\{0\}$.
@@
Let $R_{\lambda}=\mathbb{F}_{q}[x]/\langle x^{m}-\lambda\rangle$. Note that $\lambda$-constacyclic codes over $\mathbb{F}_{q}$ of length $m$ are ideals of $R_{\lambda}$. Let $C$ be a $\lambda$-quasi-twisted code of length $n=m\ell$ and index $\ell$ over $\mathbb{F}_{q}$. It is well known that $C$ can be identified as an $R_{\lambda}$-submodule of $R_{\lambda}^{\ell}$ (for instance, see [14]). The correspondence is given by
@@
Throughout the article, we assume $\gcd(m,q)=1$. We decompose quasi-cyclic codes ($\lambda=1$) into constituent codes (for details, see [13, Chapter 7, Quasi-cyclic codes p.132]). Let $x^{m}-1=f_{1}(x)f_{2}(x)\dots f_{t}(x)$ be the factorization into a product of distinct irreducible polynomials over $\mathbb{F}_{q}$. For each $i$, let $F_{i}:=\mathbb{F}_{q}[x]/\langle f_{i}(x)\rangle$. Let $\xi$ be a primitive $m^{\text{th}}$ root of unity in some extension of $\mathbb{F}_{q}$. Let $\xi^{u_{i}}$ be a root of $f_{i}(x)$. Then $F_{i}=\mathbb{F}_{q}(\xi^{u_{i}})$. By the Chinese Remainder Theorem (CRT), $R=\mathbb{F}_{q}[x]/\langle x^{m}-1\rangle$ can be decomposed as
@@
We have the following characterization of QC LCP of codes from [9, Theorem 3.1].
@@
3 LCPs of quasi-cyclic codes of index $2$
@@
Let $C$ be a quasi-cyclic code of length $2m$ and index $2$ Then $C$ is generated by two elements $(g_{11}(x),g_{12}(x))$ and $(0,g_{22}(x))$ such that they satisfy the following conditions:
@@
Moreover, in this case $\dim C=2m-\deg g_{11}(x)-\deg g_{22}(x)$.
@@
Let $g(x)=\gcd(g_{11}(x),g_{22}(x))$ and $f(x)=\gcd(f_{11}(x),f_{22}(x))$. Then, by Remark 3.2, $g(x)\mid g_{12}(x)$ and $f(x)\mid f_{12}(x)$. Let $l(x)=(x^{m}-1)/\text{lcm}(g_{11}(x),g_{22}(x))$ and $h(x)=(x^{m}-1)/\text{lcm}(f_{11}(x),f_{22}(x))$. Let $g_{11}(x)=g(x)g_{11}^{\prime}(x),g_{22}(x)=g(x)g_{22}^{\prime}(x)$, and $f_{11}(x)=f(x)f_{11}^{\prime}(x),f_{22}(x)=f(x)f_{22}^{\prime}(x)$. Then
$x^{m}-1=g(x)g_{11}^{\prime}(x)g_{22}^{\prime}(x)l(x)=f(x)f_{11}^{\prime}(x)f_{22}^{\prime}(x)h(x).$
Next, we give a new characterization for a pair of quasi-cyclic codes of index $2$ to be an LCP of codes. We use the notation as above.
@@
Let $C$ and $D$ be two quasi-cyclic codes of length $2m$ and index $2$ generated by $(g_{11}(x),g_{12}(x))$, $(0,g_{22}(x))$ and $(f_{11}(x),f_{12}(x))$, $(0,f_{22}(x))$, respectively, satisfying ()($\ast$)( ∗ ). Then $(C,D)$ is an LCP of codes if and only if all of the following conditions are true:
@@
By Lemma 3.4, $(I)$ holds. Assume $g(x)\neq\frac{x^{m}-1}{\text{lcm}(f_{11}(x),f_{22}(x))}$. This implies there exist irreducible $f_{i}(x)$ or $f_{j}(x)$ such that
$(i)$ $f_{i}(x)\mid g(x)$ and $f_{i}(x)\nmid\frac{x^{m}-1}{\text{lcm}(f_{11}(x),f_{22}(x))}$,
$(ii)$ $f_{j}(x)\nmid g(x)$ and $f_{j}(x)\mid\frac{x^{m}-1}{\text{lcm}(f_{11}(x),f_{22}(x))}$.
@@
$(ii)$ If $f_{j}(x)\nmid g(x)$ and $f_{j}(x)\mid\frac{x^{m}-1}{\text{lcm}(f_{11}(x),f_{22}(x))}$, then $f_{j}(x)\nmid g_{11}(x)$ or $f_{j}(x)\nmid g_{22}(x)$, and $f_{j}(x)\nmid\text{lcm}(f_{11}(x),f_{22}(x))$ (equivalently, $f_{j}(x)\nmid f_{11}(x),f_{22}(x)$). This implies that $g_{11}(\xi^{u_{j}})\neq 0$ or $g_{22}(\xi^{u_{j}})\neq 0$, and $f_{11}(\xi^{u_{j}})\neq 0\neq f_{22}(\xi^{u_{j}})$. Then $C_{j}$ and $D_{j}$ are generated by the rows of the matrix
$G_{j}=\begin{bmatrix}g_{11}(\xi^{u_{j}})&g_{12}(\xi^{u_{j}})\\ 0&g_{22}(\xi^{u_{j}})\end{bmatrix}\text{ and }H_{j}=\begin{bmatrix}f_{11}(\xi^{u_{j}})&f_{12}(\xi^{u_{j}})\\ 0&f_{22}(\xi^{u_{j}})\end{bmatrix},$
respectively. This implies that $\dim(C_{j})\geq 1$ and $\dim(D_{j})=2$. Thus $C_{j}\cap D_{j}\neq\{0\}$. Therefore, $(C_{j},D_{j})$ is not an LCP of codes. ∎
@@
The proof is similar to that of Lemma 3.5
Lemma 3.7.
If $(C,D)$ is an LCP of codes, then $(IV)$ holds.
Proof.
@@
[…]
@@
If $(I),(II),(III),(IV)$ hold, then $(C,D)$ is an LCP of codes, that is $C_{i}\cap D_{i}=\{0\}$ and $\dim(C_{i})+\dim(D_{i})=2$ for all $1\leq i\leq t$.
@@
Let $l(x)=(x^{m}-1)/\text{lcm}(g_{11}(x),g_{22}(x))$ and $h(x)=(x^{m}-1)/\text{lcm}(f_{11}(x),f_{22}(x))$. By $(II)$ and $(III)$, we get
$x^{m}-1=g(x)g_{11}^{\prime}(x)g_{22}^{\prime}(x)f(x)=g(x)f_{11}^{\prime}(x)f_{22}^{\prime}(x)f(x).$
For each $i$, the irreducible polynomial $f_{i}$ divides exactly one of the factors $g(x)$, $g_{11}^{\prime}(x)$, $g_{22}^{\prime}(x)$, $f(x)$ of $x^{m}-1$, which leads to the following four cases.
$(i)$ $f_{i}(x)\mid g(x)$. Then $f_{i}(x)\mid g_{11}(x),g_{22}(x),g_{12}(x)$ and $f_{i}(x)\nmid f_{11}^{\prime}(x),f_{22}^{\prime}(x),f(x)$. It follows that
@@
$(ii)$ $f_{i}(x)\mid g_{11}^{\prime}(x)$. This implies that $f_{i}(x)\nmid g(x),g_{22}^{\prime}(x),f(x)$ and then either $f_{i}(x)\mid f_{11}^{\prime}(x)$ or $f_{i}(x)\mid f_{22}^{\prime}(x)$. By $(I)$, $f_{i}(x)\nmid f_{11}(x)$ and thus $f_{i}(x)\mid f_{22}^{\prime}(x)$. It follows that
@@
$(iii)$ $f_{i}(x)\mid g_{22}^{\prime}(x)$. This implies that $f_{i}(x)\nmid g(x),g_{11}^{\prime}(x),f(x)$ and thus either $f_{i}(x)\mid f_{11}^{\prime}(x)$ or $f_{i}(x)\mid f_{22}^{\prime}(x)$.
@@ 2. If $f_{i}(x)\mid f_{22}^{\prime}(x)$ then $f_{i}(x)\nmid f_{11}^{\prime}(x)$ and thus
By $(IV)$, $f_{i}(x)\nmid(g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x))$ (since $f_{i}(x)\mid g_{22}^{\prime}(x),f_{22}^{\prime}(x)$). This implies that $g_{11}(\xi^{u_{i}})f_{12}(\xi^{u_{i}})-g_{12}(\xi^{u_{i}})f_{11}(\xi^{u_{i}})\neq 0$. Therefore, $C_{i}\cap D_{i}=\{0\}$ and $\dim(C_{i})+\dim(D_{i})=2$.
$(iv)$ $f_{i}(x)\mid f(x)$. This implies that, $f_{i}(x)\nmid g(x),g_{11}^{\prime}(x),g_{22}^{\prime}(x)$ and $f_{i}(x)\mid f_{11}(x),f_{22}(x),f_{12}(x)$. Thus
$G_{i}=\begin{bmatrix}g_{11}(\xi^{u_{i}})&g_{12}(\xi^{u_{i}})\\ 0&g_{22}(\xi^{u_{i}})\end{bmatrix},\text{ and }H_{i}=\begin{bmatrix}0&0\\ 0&0\end{bmatrix},$
where $g_{11}(\xi^{u_{i}})\neq 0\neq g_{22}(\xi^{u_{i}})$. It can then be readily checked that $C_{i}\cap D_{i}=\{0\}$ and $\dim(C_{i})+\dim(D_{i})=2$.
@@
We recall that we want to prove that $(C,D)$ is an LCP of codes if and only if the conditions (I), (II), (III), and (IV) hold. The “if” direction follows from Lemmata 3.8. The “only if” direction follows from Lemmata 3.4, 3.5, 3.6, and 3.7.
@@ 4 LCPs of one-generator quasi-cyclic codes of index 2
In this section, we consider one-generator quasi-cyclic codes of index $2$. The following lemma from [3, Lemma 2.4] shows how one-generator quasi-cyclic codes are generated by two generators satisfying ()($\ast$)( ∗ ).
@@
Then the code $C$ is generated by two elements $(g_{11}(x),g_{12}(x)\mod{g_{22}(x)})$ and $(0,g_{22}(x))$ satisfying ()($\ast$)( ∗ ) with $\gcd(g_{11}(x),g_{22}(x))=g(x)$.
From the above Lemma 4.1, if $C$ is one-generator quasi-cyclic code of index $2$, then we can assume that $C$ is generated by two elements $(g_{11}(x),\tilde{g}_{12}(x))$ and $(0,g_{22}(x))$ satisfying ()($\ast$)( ∗ ), where $\tilde{g}_{12}(x)=g_{12}(x)\mod{g_{22}(x)}$. Furthermore, also by Lemma 4.1 we have that $\gcd(g_{11}(x),g_{22}(x))=g(x)$. In addition, $(0,g_{22}(x))\in C$ (see proof of [3, Lemma 2.4]). It follows that the code $C$ generated by the element $(g_{11}(x),g_{12}(x))$ is the same as the code generated by the element $(g_{11}(x),\tilde{g}_{12}(x))$. Thus, without loss of generality, from now on, we will write $g_{12}(x)$ instead of $\tilde{g}_{12}(x)$. We have
@@
Let $g_{11}(x)$, $g_{12}(x)$, $f_{11}(x)$, and $f_{12}(x)$ in $\mathbb{F}_{q}[x]$ be such that $g_{11}(x)$ and $f_{11}(x)$ divide $x^{m}-1$. Let $\gcd(g_{11}(x),g_{12}(x))=1$, $\gcd(f_{11}(x),f_{12}(x))=1$ and $\gcd(f_{11}(x),g_{11}(x))=1$. Then the following are equivalent
@@
    $\gcd\left(\frac{x^{m}-1}{f_{11}(x)},\frac{x^{m}-1}{g_{11}(x)},g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x)\right)=1$
@@
Let $C$ and $D$ be one-generator quasi-cyclic codes generated by $(g_{11}(x),g_{12}(x))$, and $(f_{11}(x),f_{12}(x))$, respectively, where $g_{11}(x),f_{11}(x)\mid(x^{m}-1)$. Then $(C,D)$ is an LCP of codes if and only if all the following conditions hold:
  1. (A)
    $\gcd(g_{11}(x),g_{12}(x))=1$.
  2. (B)
    $\gcd(f_{11}(x),f_{12}(x))=1$.
  3. (C)
    $\gcd(x^{m}-1,g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x)=1$.
@@
Let $g(x)=\gcd(g_{11}(x),g_{12}(x))$, $g_{11}(x)=g(x)g_{11}^{\prime}(x)$, $f(x)=\gcd(f_{11}(x),f_{12}(x))$ and $f_{11}(x)=f(x)f_{11}^{\prime}(x)$. Then by Lemma 4.1, we can assume that $C$ is generated by two elements $(g_{11}(x),g_{12}(x))$ and $(0,g_{22}(x))$, and $D$ is generated by two elements $(f_{11}(x),f_{12}(x))$ and $(0,f_{22}(x))$, where
@@
It is easy to see the generating elements for $C$ and $D$ satisfy the conditions ()($\ast$)( ∗ ) and $x^{m}-1=\text{lcm}(g_{11}(x),g_{22}(x))=\text{lcm}(f_{11}(x),f_{22}(x))$ (see Eq. 4.1). By Theorem 3.3, $(C,D)$ is an LCP of codes if and only if

### v2

@@ Abstract
In this paper, we provide a polynomial characterization of linear complementary pairs of quasi-cyclic and quasi-twisted codes of index $2$. We also give several examples of linear complementary pairs of quasi-cyclic and quasi-twisted codes with optimal security parameters.
@@ 1 Introduction
In 2018, Carlet et al. [10] characterized LCPs of quasi-cyclic codes using the Chinese Remainder Theorem (CRT) decomposition of codes and also provided necessary and sufficient conditions for pairs of constacyclic and double circulant codes to be an LCP of codes. Moreover, they proved that if $(C,D)$ is an LCP of constacyclic or 2D cyclic code, then $C$ and $D^{\perp}$ are permutation equivalent. Güneri et al. [16] showed the equivalence of $C$ and $D^{\perp}$ holds for an LCP of $mD$-cyclic codes as well. These results show that finding LCD codes and LCPs of codes is the same in the context of security parameters. However, $C$ and $D^{\perp}$ need not be permutation equivalent for an LCP of quasi-cyclic codes [10, Example 3.3]. In [11], Carlet et al. proved that for $q>2$, there exists an LCP of linear codes with security parameters as good as the optimal distance of linear codes, and for $q=2$, the security parameters of optimal binary LCPs of linear codes are lower bounded by one less than the optimal distance of linear codes. In [13, 15, 21], the authors discussed optimal binary LCPs of linear codes. In this work, we give a polynomial characterization of LCPs of quasi-cyclic and quasi-twisted codes with index $2$. Using these characterizations, we provide several examples of optimal LCPs of quasi-twisted codes over $\mathbb{F}_{q}$. The rest of the paper is organized as follows.
@@ 2 Preliminaries
Note that if $C$ has parameters $[n,k]$, then $D$ must have parameters [n,nk][n,n-k]. Linear complementary dual (LCD) codes are a special class of LCPs of codes where $D$ is the dual code of $C$. Equivalently, $C$ is an LCD code if $C\cap C^{\perp}=\{0\}$.
@@
Let $R_{\lambda}=\mathbb{F}_{q}[x]/\langle x^{m}-\lambda\rangle$ and for $\lambda=1$, we denote $R_{1}$ by $R$. Note that $\lambda$-constacyclic codes over $\mathbb{F}_{q}$ of length $m$ are ideals of $R_{\lambda}$. Let $C$ be a $\lambda$-quasi-twisted code of length $n=m\ell$ and index $\ell$ over $\mathbb{F}_{q}$. It is well known that $C$ can be identified as an $R_{\lambda}$-submodule of $R_{\lambda}^{\ell}$ (for instance, see [18]). The correspondence is given by
@@
Throughout the article, we assume $\gcd(m,q)=1$. We decompose quasi-cyclic codes ($\lambda=1$) into constituent codes (for details, see [17, Chapter 7, Quasi-cyclic codes p.132]). Let $x^{m}-1=f_{1}(x)f_{2}(x)\dots f_{t}(x)$ be the factorization into a product of distinct irreducible polynomials over $\mathbb{F}_{q}$. For each $i$, let $F_{i}:=\mathbb{F}_{q}[x]/\langle f_{i}(x)\rangle$. Let $\xi$ be a primitive $m^{\text{th}}$ root of unity in some extension of $\mathbb{F}_{q}$. Let $\xi^{u_{i}}$ be a root of $f_{i}(x)$. Then $F_{i}=\mathbb{F}_{q}(\xi^{u_{i}})$. By the Chinese Remainder Theorem (CRT), $R=\mathbb{F}_{q}[x]/\langle x^{m}-1\rangle$ can be decomposed as
@@
We have the following characterization of QC LCP of codes from [10, Theorem 3.1].
@@
3 LCPs of quasi-cyclic codes of index 22
@@
Let $C$ be a quasi-cyclic code of length $2m$ and index $2.$ Then $C$ is generated by two elements $(g_{11}(x),g_{12}(x))$ and $(0,g_{22}(x))$ such that they satisfy the following conditions:
@@
Moreover, in this case $\dim(C)=2m-\deg g_{11}(x)-\deg g_{22}(x)$.
@@
Let $g(x)=\gcd(g_{11}(x),g_{22}(x))$ and $f(x)=\gcd(f_{11}(x),f_{22}(x))$. Then, by Remark 3.1, $g(x)\mid g_{12}(x)$ and $f(x)\mid f_{12}(x)$. Let $g_{11}(x)=g(x)g_{11}^{\prime}(x),g_{22}(x)=g(x)g_{22}^{\prime}(x)$, and $f_{11}(x)=f(x)f_{11}^{\prime}(x),f_{22}(x)=f(x)f_{22}^{\prime}(x)$.
Now, we give a new characterization for a pair of quasi-cyclic codes of index $2$ to be an LCP of codes. We use the notations as above.
@@
Let $C$ and $D$ be two quasi-cyclic codes of length $2m$ and index $2$ generated by $(g_{11}(x),g_{12}(x))$, $(0,g_{22}(x))$ and $(f_{11}(x),f_{12}(x))$, $(0,f_{22}(x))$, respectively, satisfying Condition $(\ast)$. Then $(C,D)$ is an LCP of codes if and only if all of the following conditions are true:
@@
Let $(C,D)$ be an LCP of codes. By Lemma 3.3, $(I)$ holds. We claim that
$g(x)=\frac{x^{m}-1}{\text{lcm}(f_{11}(x),f_{22}(x))},\text{ where }g(x)=\gcd(g_{11}(x),g_{22}(x)).$ (3.1)
On the contrary, assume $g(x)\neq\frac{x^{m}-1}{\text{lcm}(f_{11}(x),f_{22}(x))}$. This implies there exists irreducible $f_{i}(x)$ such that:
$(i)$ $f_{i}(x)\mid g(x)$ and $f_{i}(x)\nmid\frac{x^{m}-1}{\text{lcm}(f_{11}(x),f_{22}(x))}$ or
$(ii)$ $f_{i}(x)\nmid g(x)$ and $f_{i}(x)\mid\frac{x^{m}-1}{\text{lcm}(f_{11}(x),f_{22}(x))}$.
@@
$(ii)$ If $f_{i}(x)\nmid g(x)$ and $f_{i}(x)\mid\frac{x^{m}-1}{\text{lcm}(f_{11}(x),f_{22}(x))}$, then $f_{i}(x)\nmid g_{11}(x)$ or $f_{i}(x)\nmid g_{22}(x)$, and $f_{i}(x)\nmid\text{lcm}(f_{11}(x),f_{22}(x))$ (equivalently, $f_{i}(x)\nmid f_{11}(x),f_{22}(x)$). This implies that $g_{11}(\xi^{u_{i}})\neq 0$ or $g_{22}(\xi^{u_{i}})\neq 0$, and $f_{11}(\xi^{u_{i}})\neq 0\neq f_{22}(\xi^{u_{i}})$. Then $C_{i}$ and $D_{i}$ are generated by the rows of the matrix
respectively. This implies that $\dim(C_{i})\geq 1$ and $\dim(D_{i})=2$. Thus $C_{i}\cap D_{i}\neq\{0\}$. Therefore, $(C_{i},D_{i})$ is not an LCP of codes, a contradiction. Thus, Eq. 3.1 holds.
Using similar arguments, we get
$f(x)=\frac{x^{m}-1}{\text{lcm}(g_{11}(x),g_{22}(x))},\text{ where }f(x)=\gcd(f_{11}(x),f_{22}(x)).$ (3.2)
By Eqs. 3.1 and 3.2, we have
$g(x)\text{lcm}(f_{11}(x),f_{22}(x))\cdot f(x)\text{lcm}(g_{11}(x),g_{22}(x))=(x^{m}-1)^{2},$
$i.e.,\ \$$g_{11}(x)g_{22}(x)f_{11}(x)f_{22}(x)=(x^{m}-1)^{2}.$
This completes the proof. ∎
@@
 $\text{ and }H_{i}=\begin{bmatrix}f_{11}(\xi^{u_{i}})&f_{12}(\xi^{u_{i}})\\ 0&f_{22}(\xi^{u_{i}})\end{bmatrix}=\begin{bmatrix}f_{11}(\xi^{u_{i}})&f_{12}(\xi^{u_{i}})\\ 0&0\end{bmatrix}$,
@@
If $(I),(II),(III)$ hold, then $(C,D)$ is an LCP of codes, that is $C_{i}\cap D_{i}=\{0\}$ and $\dim(C_{i})+\dim(D_{i})=2$ for all $1\leq i\leq t$.
@@
By $(II)$, we have
$g(x)g(x)g_{11}^{\prime}(x)g_{22}^{\prime}(x)f_{11}^{\prime}(x)f_{22}^{\prime}(x)f(x)f(x)=(x^{m}-1)^{2},$ (3.3)
where $g(x)=\gcd(g_{11}(x),g_{22}(x))$ and $f(x)=\gcd(f_{11}(x),f_{22}(x))$. For each $i$, the irreducible polynomial $f_{i}$ divides at least one of the factors $g(x)$, $g_{11}^{\prime}(x)$, $g_{22}^{\prime}(x)$, $f(x),f_{11}^{\prime}(x),f_{22}^{\prime}(x)$ of $x^{m}-1$, which leads to the following six cases.
$\bm{(i)}$ Let $f_{i}(x)\mid g(x)$. Then $f_{i}(x)\mid g_{11}(x),g_{22}(x),g_{12}(x)$. Also, $f_{i}^{3}(x)\nmid(x^{m}-1)^{2}$ implies that $f_{i}(x)\nmid f_{11}^{\prime}(x),f_{22}^{\prime}(x),f(x)$. It follows that
@@
$\bm{(ii)}$ Let $f_{i}(x)\mid g_{11}^{\prime}(x)$. Then $f_{i}(x)\nmid g_{22}^{\prime}(x)$. Also, $f_{i}^{3}(x)\nmid(x^{m}-1)^{2}$ implies that $f_{i}(x)\nmid g(x),f(x)$. Hence, either $f_{i}(x)\mid f_{11}^{\prime}(x)$ or $f_{i}(x)\mid f_{22}^{\prime}(x)$. By $(I)$, $f_{i}(x)\nmid f_{11}(x)$ and thus $f_{i}(x)\mid f_{22}^{\prime}(x)$. It follows that
@@
$\bm{(iii)}$ Let $f_{i}(x)\mid g_{22}^{\prime}(x)$. Then $f_{i}(x)\nmid g(x),g_{11}^{\prime}(x),f(x)$. Thus either $f_{i}(x)\mid f_{11}^{\prime}(x)$ or $f_{i}(x)\mid f_{22}^{\prime}(x)$.
@@ 2. If $f_{i}(x)\mid f_{22}^{\prime}(x)$ then $f_{i}(x)\nmid f_{11}^{\prime}(x)$ and thus
By $(III)$, $f_{i}(x)\nmid(g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x))$ (since $f_{i}(x)\mid g_{22}^{\prime}(x),f_{22}^{\prime}(x)$). This implies that $g_{11}(\xi^{u_{i}})f_{12}(\xi^{u_{i}})-g_{12}(\xi^{u_{i}})f_{11}(\xi^{u_{i}})\neq 0$. Therefore, $C_{i}\cap D_{i}=\{0\}$ and $\dim(C_{i})+\dim(D_{i})=2$.
$\bm{(iv)}$ Let $f_{i}(x)\mid f(x)$. The proof is similar to the case $\bm{(i)}$ by interchanging codes $C_{i}$ and $D_{i}$.
$\bm{(v)}$ Let $f_{i}(x)\mid f_{11}^{\prime}(x)$. The proof is similar to the case $\bm{(ii)}$ by interchanging codes $C_{i}$ and $D_{i}$.
$\bm{(vi)}$ Let $f_{i}(x)\mid f_{22}^{\prime}(x)$. The proof is similar to the case $\bm{(iii)}$ by interchanging codes $C_{i}$ and $D_{i}$. ∎
@@
We recall that we want to prove that $(C,D)$ is an LCP of codes if and only if the conditions (I), (II), and (III) hold. The “if” direction follows from Lemma 3.6. The “only if” direction follows from Lemmata 3.3, 3.4, and 3.5.
@@ 4 LCPs of one-generator quasi-cyclic codes of index 2
In this section, we consider one-generator quasi-cyclic codes of index $2$. The following lemma from [3, Lemma 2.4] shows how one-generator quasi-cyclic codes are generated by two generators satisfying Condition $(\ast)$.
@@
Then the code $C$ is generated by two elements $(g_{11}(x),g_{12}(x)\pmod{g_{22}(x)})$ and $(0,g_{22}(x))$ satisfying Condition $(\ast)$ with $\gcd(g_{11}(x),g_{22}(x))=g(x)$.
From the above Lemma 4.1, if $C$ is one-generator quasi-cyclic code of index $2$, then we can assume that $C$ is generated by two elements $(g_{11}(x),\tilde{g}_{12}(x))$ and $(0,g_{22}(x))$ satisfying Condition $(\ast)$, where $\tilde{g}_{12}(x)=g_{12}(x)\pmod{g_{22}(x)}$. Furthermore, also by Lemma 4.1 we have that $\gcd(g_{11}(x),g_{22}(x))=g(x)$. In addition, $(0,g_{22}(x))\in C$ (see proof of [3, Lemma 2.4]). It follows that code $C$ generated by the element $(g_{11}(x),g_{12}(x))$ is the same as the code generated by the element $(g_{11}(x),\tilde{g}_{12}(x))$. Thus, without loss of generality, from now on, we will write $g_{12}(x)$ instead of $\tilde{g}_{12}(x)$. We have
@@
Let $g_{11}(x)$, $g_{12}(x)$, $f_{11}(x)$, and $f_{12}(x)$ in $\mathbb{F}_{q}[x]$ be such that $g_{11}(x)$ and $f_{11}(x)$ divide $x^{m}-1$. Let $\gcd(g_{11}(x),g_{12}(x))=1$, $\gcd(f_{11}(x),f_{12}(x))=1$ and $\gcd(f_{11}(x),g_{11}(x))=1$. Then the following are equivalent.
@@
    $\gcd\left(\frac{x^{m}-1}{f_{11}(x)},\frac{x^{m}-1}{g_{11}(x)},g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x)\right)=1$.
@@
Let $C$ and $D$ be one-generator quasi-cyclic codes generated by $(g_{11}(x),g_{12}(x))$ and $(f_{11}(x),f_{12}(x))$, respectively, where $g_{11}(x),f_{11}(x)\mid(x^{m}-1)$. Then $(C,D)$ is an LCP of codes if and only if $\gcd(x^{m}-1,g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x))=1$.
@@
Let $g(x)=\gcd(g_{11}(x),g_{12}(x))$, $g_{11}(x)=g(x)g_{11}^{\prime}(x)$, $f(x)=\gcd(f_{11}(x),f_{12}(x))$, and $f_{11}(x)=f(x)f_{11}^{\prime}(x)$. Then by Lemma 4.1, we can assume that $C$ is generated by two elements $(g_{11}(x),g_{12}(x))$ and $(0,g_{22}(x))$, and $D$ is generated by two elements $(f_{11}(x),f_{12}(x))$ and $(0,f_{22}(x))$, where
@@
It is easy to see the generating elements for $C$ and $D$ satisfy Condition $(\ast)$ and $x^{m}-1=\text{lcm}(g_{11}(x),g_{22}(x))=\text{lcm}(f_{11}(x),f_{22}(x))$ (see Eq. 4.1). By Theorem 3.2, $(C,D)$ is an LCP of codes if and only if